1. What is the present value of a $500,000 inheritance (trust fund) to be received in 20 years discounted at an annual rate of 12%?
Answer #1

FV =500000
PV =FV/(1+r)^n =500000/(1+12%)^20 =51833.38
PV of trust fund =51833.38